I need to have four different rectangles whose perimeter equals 6 and a half. Can you help me out?
Answer by richwmiller(17219)

2*(W+L)=6.5
(W+L)=6.5/2
W=1/2*(6.5-2L)
Pick numbers for L and solve for W
L=1 then W=4.5/2=2.25
L=2 then W=2.5/2=1.25
L=3 then W=.5/2=.25
L=4 then W=-1.5/2=-.75
L can't be 4
Let L=2.5 then W=1.5/2=.75